### v1

# On three recent questions of Bourin and Lee on quadratic symmetric modulus and Euler-operator identity

Abstract.
We answer three questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [6]. First, we show that the exponent 22 in their unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by an explicit $2\times 2$ counterexample for $p>2$. Second, we construct a compact operator for which the corresponding singular-value inequality fails for every $p>2$ (indeed for all $p>0$ at a fixed pair of indices). Finally, we obtain isometry-orbit refinements of Euler’s quadrilateral identity for matrices and derive several Clarkson–McCarthy type inequalities for Schatten $p$-norms and related consequences.
Key words and phrases:
Symmetric modulus; singular values; unitary orbits; Euler identity; Schatten norms
2020 Mathematics Subject Classification:
15A60, 47A30, 47A57, 15A42
1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space, and let $\mathbb{K}$ denote the ideal of compact operators on $\mathcal{H}$. For $Z\in\mathbb{K}$, the singular values $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ are defined as the eigenvalues (counted with multiplicity) of $|Z|=(Z^{*}Z)^{1/2}$, listed in nonincreasing order. Let $\mathbb{M}_{m,n}$ be the set of all complex $m\times n$ matrices; when $m=n$, we write $\mathbb{M}_{n}$. For a matrix $X\in\mathbb{M}_{n}$, we adopt the compatible convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ may be viewed as an infinite nonincreasing sequence (and this agrees with the compact-operator notation).
We write
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2},\qquad\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}.$
For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite.
For $p>0$, the Schatten pp-(quasi)norm on $\mathbb{M}_{n}$ is
$\|X\|_{p}:=\big(\operatorname{Tr}|X|^{p}\big)^{1/p},\qquad X\in\mathbb{M}_{n}.$
When $1<p<\infty$, we denote by $q$ the conjugate exponent, i.e. $1/p+1/q=1$.
Bourin and Lee [6, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus.
Theorem 1.1 (Bourin–Lee).
Let $Z\in\mathbb{M}_{n}$. Then there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that
$\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\leq U|\operatorname{Re}\,Z|U^{*}+V|\operatorname{Im}\,Z|V^{*}.$
They also posed the following question concerning the optimality of the exponent 22.
Question 1.2 ([6, Question 4.5]).
Let $p>0$. Suppose that, for every $Z\in\mathbb{M}_{n}$, one can find unitary matrices $U,V\in\mathbb{M}_{n}$ such that
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq U\,|\operatorname{Re}\,Z|\,U^{*}+V\,|\operatorname{Im}\,Z|\,V^{*}.$ (1.1)
Must we necessarily have $p\leq 2$?
Our first result answers Question 1.2 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
Theorem 1.3.
Let $n\geq 2$ and $p>0$. If for every $Z\in\mathbb{M}_{n}$ there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that (1.1) holds, then necessarily $p\leq 2$.
Remark 1.4.
For $n=1$, the statement of Question 1.2 is different: for a complex number $z$, the left-hand side of (1.1) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$. Hence (1.1) holds for all $p>0$ in dimension 11. The necessity $p\leq 2$ is therefore a genuinely noncommutative phenomenon requiring $n\geq 2$.
Bourin and Lee [6] further raised the following question on singular values.
Question 1.5 ([6, Question 4.7]).
Does there exist $Z\in\mathbb{K}$ such that for every $p>2$ one has
$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j+k}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
for some pair of integers $j,k\geq 0$?
We answer this question in the affirmative; in fact, we produce an example where the same pair of indices works for all $p>0$.
Theorem 1.6.
There exists $Z\in\mathbb{K}$ such that for every $p>2$ one can find integers $j,k\geq 0$ for which
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z).$ (1.2)
Moreover, one may take $j=k=1$, in which case the inequality holds for all $p>0$.
Finally, Bourin and Lee [6] asked for matrix analogues of Euler’s quadrilateral identity.
Question 1.7 ([6, Question 4.8]).
What are the matrix versions, if any, of Euler’s quadrilateral identity
$\|x+y+z\|^{2}+\|x\|^{2}+\|y\|^{2}+\|z\|^{2}=\|x+y\|^{2}+\|y+z\|^{2}+\|z+x\|^{2}$
for three points $x,y,z\in\mathbb{C}^{n}$?
A direct computation yields the following Euler-type operator identity.
Proposition 1.8 (Euler operator identity).
Let $A,B,C\in\mathbb{M}_{n}$. Then
$|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}=|A+B|^{2}+|B+C|^{2}+|C+A|^{2}.$ (1.3)
We recall that an isometry is a rectangular matrix $W\in\mathbb{M}_{m,n}$ such that $W^{*}W=I_{n}$. Bourin and Lee [6] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following main results give isometry-orbit refinements of (1.3).
Theorem 1.9.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist isometries $U_{ij}\in\mathbb{M}_{4n,n}$, $1\leq i,j\leq 4$, such that
$|A+B|^{2}\oplus|B+C|^{2}\oplus|C+A|^{2}\oplus 0$ $=\frac{1}{4}\sum_{j=1}^{4}U_{1j}\,|A+B+C|^{2}\,U_{1j}^{*}+\frac{1}{4}\sum_{j=1}^{4}U_{2j}\,|A|^{2}\,U_{2j}^{*}$
$\quad+\frac{1}{4}\sum_{j=1}^{4}U_{3j}\,|B|^{2}\,U_{3j}^{*}+\frac{1}{4}\sum_{j=1}^{4}U_{4j}\,|C|^{2}\,U_{4j}^{*}.$ (1.4)
Theorem 1.10.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist three isometries $U_{1},U_{2},U_{3}\in\mathbb{M}_{3n,n}$ such that
$|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}=\frac{1}{3}\sum_{k=1}^{3}U_{k}\,\left(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\right)\,U_{k}^{*}.$
Theorem 1.11.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist four isometries $U_{1},U_{2},U_{3},U_{4}\in\mathbb{M}_{4n,n}$ such that
$|A+B+C|^{2}\oplus|A|^{2}\oplus|B|^{2}\oplus|C|^{2}=\frac{1}{4}\sum_{k=1}^{4}U_{k}\,\left(|A+B|^{2}+|B+C|^{2}+|A+C|^{2}\right)\,U_{k}^{*}.$
Based on results obtained by the author in [8], we obtain the following two Clarkson–McCarthy type inequalities.
Corollary 1.12.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$,
$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}$ $\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big),$
$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}$ $\leq 3^{\frac{p}{2}-1}\big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\big).$
For $0<p\leq 2$, both inequalities reverse.
Corollary 1.13.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $1<p\leq 2$ ,
$\|A+B\|_{p}^{q}+\|B+C\|_{p}^{q}+\|C+A\|_{p}^{q}\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\Big)^{q/p},$
where $q$ is the conjugate exponent of pp. For $p\geq 2$, the inequality is reversed.
Remark 1.14.
Set C=BC=-B in Corollary 1.12 and Corollary 1.13, we derive the classic Clarkson–McCarthy inequality in [7].
Organization of this paper. In Section 2, we prove Theorem 1.3. In Section 3, we prove Theorem 1.6. In Section 4, we prove Theorems 1.9, 1.10, 1.11 and derive further related consequences.
2. Proof of Theorem 1.3
2.1. Trace necessary condition
Assume (1.1) holds for a given $Z\in\mathbb{M}_{n}$ and some unitaries $U,V$. Then, taking traces yields
$\operatorname{Tr}\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq\operatorname{Tr}|\operatorname{Re}\,Z|+\operatorname{Tr}|\operatorname{Im}\,Z|.$ (2.1)
Indeed, all terms in (1.1) are positive semidefinite, and for $0\leq A\leq B$ we have $\operatorname{Tr}A\leq\operatorname{Tr}B$. Moreover, the right-hand side of (1.1) has trace independent of the choice of $U,V$ by unitary invariance. Consequently, if (2.1) fails for some $Z$, then (1.1) fails for that $Z$ for all choices of $U,V$.
Thus, to disprove (1.1) for $p>2$, it suffices to find, for a given $p$, a matrix $Z$ violating (2.1).
$2.$ A $2\times 2$ counterexample for $p>2$
Fix $\theta\in(0,\pi/2)$ and set $c=\cos\theta$, $s=\sin\theta$. Define
$Z_{\theta}=\begin{pmatrix}c&0\\ -s&0\end{pmatrix}\in\mathbb{M}_{2}.$ (2.2)
Note that $Z_{\theta}$ is real, hence $Z_{\theta}^{*}=Z_{\theta}^{\mathsf{T}}$.
Claim 2.1.
For every $p>0$,
$\operatorname{Tr}\left(\frac{|Z_{\theta}|^{p}+|Z_{\theta}^{*}|^{p}}{2}\right)^{1/p}=\left(\frac{1+c}{2}\right)^{1/p}+\left(\frac{1-c}{2}\right)^{1/p}.$ (2.3)
Proof.
Compute
$Z_{\theta}^{*}Z_{\theta}=\begin{pmatrix}c\\ -s\end{pmatrix}(c,\,-s)=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=:P.$
Hence $|Z_{\theta}|=(Z_{\theta}^{*}Z_{\theta})^{1/2}=P$. Similarly,
$Z_{\theta}Z_{\theta}^{*}=\begin{pmatrix}c^{2}&-cs\\ -cs&s^{2}\end{pmatrix}=:Q.$
A direct computation shows $Q^{2}=Q$. Moreover $Q=Q^{*}$ (indeed $Q$ is real symmetric), hence $Q$ is an orthogonal projection. Therefore $|Z_{\theta}^{*}|=(Z_{\theta}Z_{\theta}^{*})^{1/2}=Q$.
Since $P$ and $Q$ are projections, $P^{p}=P$ and $Q^{p}=Q$ for every $p>0$. Thus
$\left(\frac{|Z_{\theta}|^{p}+|Z_{\theta}^{*}|^{p}}{2}\right)^{1/p}=\left(\frac{P+Q}{2}\right)^{1/p}.$
We compute the eigenvalues of $(P+Q)/2$. We have
$P+Q=\begin{pmatrix}1+c^{2}&-cs\\ -cs&s^{2}\end{pmatrix},\qquad\operatorname{Tr}(P+Q)=2,\qquad\det(P+Q)=s^{2}.$
Hence the characteristic polynomial is $\lambda^{2}-2\lambda+s^{2}=0$, so the eigenvalues are $\lambda_{\pm}=1\pm\sqrt{1-s^{2}}=1\pm c$. Therefore the eigenvalues of $(P+Q)/2$ are $(1\pm c)/2$, and those of $\bigl((P+Q)/2\bigr)^{1/p}$ are $\bigl((1\pm c)/2\bigr)^{1/p}$. Taking the trace yields (2.3). ∎
Claim $2.$
For $Z_{\theta}$ in (2.2),
$\operatorname{Tr}|\operatorname{Re}\,Z_{\theta}|+\operatorname{Tr}|\operatorname{Im}\,Z_{\theta}|=1+s.$ (2.4)
Proof.
We compute
$\operatorname{Re}\,Z_{\theta}=\frac{Z_{\theta}+Z_{\theta}^{*}}{2}=\begin{pmatrix}c&-s/2\\ -s/2&0\end{pmatrix},\qquad\operatorname{Im}\,Z_{\theta}=\frac{Z_{\theta}-Z_{\theta}^{*}}{2i}=\begin{pmatrix}0&-is/2\\ is/2&0\end{pmatrix}$.
For $\operatorname{Re}\,Z_{\theta}$, the characteristic polynomial is
$\lambda^{2}-c\lambda-\frac{s^{2}}{4},$
whose roots are
$\lambda_{\pm}=\frac{c\pm\sqrt{c^{2}+s^{2}}}{2}=\frac{c\pm 1}{2}.$
Since $c\in(0,1)$, we have $\lambda_{+}>0$ and $\lambda_{-}<0$, hence
$\operatorname{Tr}|\operatorname{Re}\,Z_{\theta}|=|\lambda_{+}|+|\lambda_{-}|=\frac{1+c}{2}+\frac{1-c}{2}=1.$
For $\operatorname{Im}\,Z_{\theta}$, the eigenvalues are $\pm s/2$, hence $\operatorname{Tr}|\operatorname{Im}\,Z_{\theta}|=s$. Adding yields (2.4). ∎
Combining (2.1), Claims 2.1 and 2.2, we see that if (1.1) were to hold for $Z_{\theta}$, then necessarily
$\left(\frac{1+c}{2}\right)^{1/p}+\left(\frac{1-c}{2}\right)^{1/p}\leq 1+s.$ (2.5)
Claim 2.3.
Let $p>2$. Then there exists $\theta_{0}\in(0,\pi/2)$ such that for all $\theta\in(0,\theta_{0})$ the inequality (2.5) fails.
Proof.
Define
$\Phi(\theta):=\left(\frac{1+\cos\theta}{2}\right)^{1/p}+\left(\frac{1-\cos\theta}{2}\right)^{1/p}-(1+\sin\theta).$
As $\theta\to 0^{+}$ we have the standard expansions
$\sin\theta=\theta+O(\theta^{3}),\qquad\cos\theta=1-\frac{\theta^{2}}{2}+O(\theta^{4}).$
Hence
$\frac{1+\cos\theta}{2}=1-\frac{\theta^{2}}{4}+O(\theta^{4})\quad\Longrightarrow\quad\left(\frac{1+\cos\theta}{2}\right)^{1/p}=1+O(\theta^{2}),$
and
$\frac{1-\cos\theta}{2}=\frac{\theta^{2}}{4}+O(\theta^{4})=\frac{\theta^{2}}{4}\bigl(1+O(\theta^{2})\bigr)$
$\Longrightarrow\left(\frac{1-\cos\theta}{2}\right)^{1/p}=4^{-1/p}\theta^{2/p}\bigl(1+O(\theta^{2})\bigr)=4^{-1/p}\theta^{2/p}+O(\theta^{2/p+2}).$
Combining these estimates gives
$\Phi(\theta)=4^{-1/p}\theta^{2/p}-\theta+O(\theta^{2})+O(\theta^{2/p+2})\qquad(\theta\to 0^{+}).$
Dividing by $\theta$ yields
$\frac{\Phi(\theta)}{\theta}=4^{-1/p}\theta^{2/p-1}-1+O(\theta)+O(\theta^{2/p+1})\xrightarrow[\theta\to 0^{+}$$\infty$,
since $p>2$ implies 2/p1<02/p-1<0. Therefore $\Phi(\theta)>0$ for all sufficiently small $\theta$, i.e. (2.5) fails for all $\theta\in(0,\theta_{0})$ for some $\theta_{0}>0$. ∎
Proof of Theorem 1.3.
Assume $p>2$. Consider the $2\times 2$ matrix $Z_{\theta}$ defined in (2.2). If (1.1) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (2.1). Claims 2.1 and 2.2 reduce (2.1) to (2.5), which fails for all sufficiently small $\theta$ by Claim 2.3. Hence (1.1) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
If $n>2$, embed $Z_{\theta}$ as $Z_{\theta}\oplus 0_{n-2}\in\mathbb{M}_{n}$; the same trace obstruction applies, so (1.1) fails in $\mathbb{M}_{n}$ as well. This contradicts the hypothesis, so necessarily $p\leq 2$. ∎
3. Proof of Theorem 1.6
Proof of Theorem 1.6.
Let $H=\ell^{2}(\mathbb{N})$ with its standard orthonormal basis $(e_{n})_{n\geq 1}$, and define
$Ze_{1}=e_{1},\qquad Ze_{2}=e_{2},\qquad Ze_{3}=i\,e_{3},\qquad Ze_{n}=0\ \ (n\geq 4).$
Equivalently,
$Z=\operatorname{diag}(1,\,1,\,i,\,0,\,0,\dots).$
This is a rank-33 operator, hence $Z\in\mathbb{K}$.
Since $Z$ is diagonal (hence normal), we have $|Z|=|Z^{*}|$. Indeed,
$Z^{*}=\operatorname{diag}(1,\,1,\,-i,\,0,\,0,\dots),\qquad Z^{*}Z=\operatorname{diag}(1,\,1,\,1,\,0,\,0,\dots),$
and therefore
[…]
Consequently, for every $p>0$,
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}=|Z|.$
Hence the singular values of the left-hand operator are
$\mu_{1}^{\downarrow}=1,\ \mu_{2}^{\downarrow}=1,\ \mu_{3}^{\downarrow}=1,\ \mu_{m}^{\downarrow}=0\ (m\geq 4).$
Next,
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2}=\operatorname{diag}(1,\,1,\,0,\,0,\,0,\dots),$
so $\mu_{3}^{\downarrow}(\operatorname{Re}Z)=0$. Also,
$\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}=\operatorname{diag}(0,\,0,\,1,\,0,\,0,\dots),$
so $\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0$.
Choose $j=k=1$, so that $1+j+k=3$ and $1+k=2$. Then for every $p>0$,
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=\mu_{3}^{\downarrow}(|Z|)=1,$
while
$\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z)=\mu_{3}^{\downarrow}(\operatorname{Re}Z)+\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0+0=0.$
Therefore (1.2) holds strictly for all $p>0$, and in particular for all $p>2$. ∎
Remark 3.1.
The construction is intentionally minimal and answers the existence question in Question 1.5. Here $Z$ is finite rank and normal, so the left-hand side of (1.2) is independent of $p$ and equals a fixed singular value of $|Z|$. The strict inequality is obtained by selecting indices ($j=k=1$) for which the corresponding singular values of $\operatorname{Re}Z$ and $\operatorname{Im}Z$ vanish, while the chosen singular value of $|Z|$ remains nonzero.
4. Euler-type identities and isometry orbits
4.1. Euler-type identities
Let
$X:=\begin{pmatrix}A+B+C\\ A\\ B\\ C\end{pmatrix},\qquad Y:=\begin{pmatrix}A+B\\ B+C\\ C+A\\ 0\end{pmatrix}\in\mathbb{M}_{4n,n}.$
Consider the $4\times 4$ Hadamard matrix
$H=\frac{1}{2}\begin{pmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ -1&1&1&-1\end{pmatrix},\qquad H^{*}H=I_{4},$
and set $\mathcal{U}:=H\otimes I_{n}\in\mathbb{M}_{4n}$. A direct block computation gives $X=\mathcal{U}\,Y$. Hence $X^{*}X=Y^{*}Y$, i.e. (1.3).
We use the following Pythagoras theorem for partitioned matrices proved by Bourin and Lee [5].
Lemma 4.1 ([5, Theorem 2.1]).
Let $m,n\geq 1$ and let $T=[T_{ij}]_{i,j=1}^{m}$ be an $m\times m$ block matrix with $T_{ij}\in\mathbb{M}_{n}$. Then there exist isometries $W_{ij}\in\mathbb{M}_{mn,n}$ such that
$|T|^{2}=\sum_{i=1}^{m}\sum_{j=1}^{m}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
Remark 4.2.
The standard $m\times m$ grid partitioning is row/column compatible in the sense of [5], so the above lemma applies directly to block matrices $T=[T_{ij}]$ written with respect to such a partition.
Next, we prove Theorem 1.9.
Proof of Theorem 1.9.
Let
$\Delta_{Y}:=\begin{pmatrix}A+B&0&0&0\\ 0&B+C&0&0\\ 0&0&C+A&0\\ 0&0&0&0\end{pmatrix}\in\mathbb{M}_{4n},\qquad\mathcal{U}:=H\otimes I_{n}.$
Set
$T:=\mathcal{U}\,\Delta_{Y}\,\mathcal{U}^{*}=\frac{1}{2}\begin{pmatrix}A+B+C&A&B&C\\ A&A+B+C&-C&-B\\ B&-C&A+B+C&-A\\ C&-B&-A&A+B+C\end{pmatrix}$.
Since $\mathcal{U}$ is unitary,
$|T|^{2}=|\mathcal{U}\Delta_{Y}\mathcal{U}^{*}|^{2}=\mathcal{U}\,|\Delta_{Y}|^{2}\,\mathcal{U}^{*}=\mathcal{U}\,\operatorname{diag}(|A+B|^{2},\ |B+C|^{2},\ |C+A|^{2},\ 0)\,\mathcal{U}^{*}.$
Apply Lemma 4.1 with $m=4$ to the $4\times 4$ block matrix $T=[T_{ij}]$. We obtain isometries $W_{ij}\in\mathbb{M}_{4n,n}$ such that
$|T|^{2}=\sum_{i=1}^{4}\sum_{j=1}^{4}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
Every block $T_{ij}$ equals $(A+B+C)/2$, $\pm A/2$, $\pm B/2$ or $\pm C/2$. Hence
$|T_{ij}|^{2}\in\left\{\tfrac{1}{4}|A+B+C|^{2},\ \tfrac{1}{4}|A|^{2},\ \tfrac{1}{4}|B|^{2},\ \tfrac{1}{4}|C|^{2}\right\}.$
Moreover, each of (A+B+C)/2,A/2,B/2,C/2(A+B+C)/2,\ A/2,\ B/2,\ C/2 appears exactly four times among the sixteen blocks. Grouping the corresponding four terms and renaming the isometries, we obtain isometries $V_{ij}\in\mathbb{M}_{4n,n}$ such that
$|T|^{2}=\frac{1}{4}\sum_{j=1}^{4}V_{1j}\,|A+B+C|^{2}\,V_{1j}^{*}+\frac{1}{4}\sum_{j=1}^{4}V_{2j}\,|A|^{2}\,V_{2j}^{*}+\frac{1}{4}\sum_{j=1}^{4}V_{3j}\,|B|^{2}\,V_{3j}^{*}+\frac{1}{4}\sum_{j=1}^{4}V_{4j}\,|C|^{2}\,V_{4j}^{*}.$
Conjugating by $\mathcal{U}^{*}$ yields
$|\Delta_{Y}|^{2}=\frac{1}{4}\sum_{j=1}^{4}(\mathcal{U}^{*}V_{1j})\,|A+B+C|^{2}\,(\mathcal{U}^{*}V_{1j})^{*}+\cdots+\frac{1}{4}\sum_{j=1}^{4}(\mathcal{U}^{*}V_{4j})\,|C|^{2}\,(\mathcal{U}^{*}V_{4j})^{*}.$
Finally, each $\mathcal{U}^{*}V_{ij}$ is still an isometry in $\mathbb{M}_{4n,n}$ since
$(\mathcal{U}^{*}V_{ij})^{*}(\mathcal{U}^{*}V_{ij})=V_{ij}^{*}\mathcal{U}\mathcal{U}^{*}V_{ij}=V_{ij}^{*}V_{ij}=I_{n}.$
Renaming $U_{ij}:=\mathcal{U}^{*}V_{ij}$ gives (1.9). ∎
Naturally, we ask the following question.
Question 4.3.
Does there exist isometries $U_{i}\in\mathbb{M}_{3n,n}$ ($1\leq i\leq 4$) such that
$|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}$ $=U_{1}\,|A+B+C|^{2}\,U_{1}^{*}+U_{2}\,|A|^{2}\,U_{2}^{*}$
$\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
The answer to Question 4.3 is negative. In fact, take $A=B=C=I_{n}$. Then
$|A+B|^{2}=|2I_{n}|^{2}=4I_{n},\qquad|A+B+C|^{2}=|3I_{n}|^{2}=9I_{n},\qquad|A|^{2}=|B|^{2}=|C|^{2}=I_{n}.$
Hence the desired identity would become
$4I_{3n}=9U_{1}U_{1}^{*}+U_{2}U_{2}^{*}+U_{3}U_{3}^{*}+U_{4}U_{4}^{*}.$
Since all terms on the right-hand side are positive semidefinite, we have
$4I_{3n}\geq 9U_{1}U_{1}^{*}.$
But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|=9$ while $\|4I_{3n}\|=4$, a contradiction. Therefore such isometries cannot exist in general.
Lemma 4.4 ([4, Lemma 3.4]).
Let
$\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$
be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
$\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}=U\begin{pmatrix}X&0\\ 0&0\end{pmatrix}U^{*}+V_{0}\begin{pmatrix}0&0\\ 0&Z\end{pmatrix}V_{0}^{*}.$
As an immediate corollary of Lemma 4.4, we obtain the following result. For completeness, we provide an alternative proof.
Lemma 4.5.
Let $m\geq 2$ and let $H=[H_{ij}]_{i,j=1}^{m}\in\mathbb{M}_{mn}$ be positive semidefinite, written in $n\times n$ blocks. Then there exist isometries $V_{1},\dots,V_{m}\in\mathbb{M}_{mn,n}$ (i.e., $V_{k}^{*}V_{k}=I_{n}$) such that
$H=\sum_{k=1}^{m}V_{k}\,H_{kk}\,V_{k}^{*}.$ (4.1)
Proof.
Since $H\geq 0$, let $R:=H^{1/2}\in\mathbb{M}_{mn}$, so that $H=R^{*}R$. Partition $R$ into $m$ block columns of width $n$:
$R=\bigl[\,R_{1}\;R_{2}\;\cdots\;R_{m}\,\bigr],\qquad R_{k}\in\mathbb{M}_{mn,n}.$
Then
$H=R^{*}R=\sum_{k=1}^{m}R_{k}R_{k}^{*}.$ (4.2)
Step 1: Identify the diagonal blocks. By block multiplication, the $(k,k)$ diagonal block of $H=R^{*}R$ is
$H_{kk}=R_{k}^{*}R_{k}\in\mathbb{M}_{n}^{+}.$
In particular, $|R_{k}|:=(R_{k}^{*}R_{k})^{1/2}=H_{kk}^{1/2}$.
Step 2: Polar decomposition and a partial isometry. Take the polar decomposition of each $R_{k}$:
$R_{k}=U_{k}\,|R_{k}|=U_{k}\,H_{kk}^{1/2},$
where $U_{k}\in\mathbb{M}_{mn,n}$ is a partial isometry satisfying
$U_{k}^{*}U_{k}=P_{k},\qquad P_{k}:=\mathrm{supp}(|R_{k}|)=\mathrm{supp}(H_{kk})$
(the orthogonal projection onto $\overline{\mathrm{Ran}}(|R_{k}|)$). Note that $H_{kk}^{1/2}(I_{n}-P_{k})=0$.
Step 3: Extend $U_{k}$ to a genuine isometry $V_{k}$. If $P_{k}=I_{n}$, then $U_{k}^{*}U_{k}=I_{n}$ and $U_{k}$ is already an isometry; set $V_{k}:=U_{k}$.
Assume now $P_{k}\neq I_{n}$. Set $d_{k}:=\mathrm{rank}(P_{k})\leq n$. Then $\dim\mathrm{Ran}(I_{n}-P_{k})=n-d_{k}$. Also, since $U_{k}$ is a partial isometry with initial projection $P_{k}$, its range projection $Q_{k}:=U_{k}U_{k}^{*}$ has rank $d_{k}$, hence
$\dim\mathrm{Ran}(I_{mn}-Q_{k})=mn-d_{k}\ \geq\ n-d_{k},$
using $m\geq 2\Rightarrow mn\geq n$.
Therefore, there exists an isometry
$W_{k}:\mathrm{Ran}(I_{n}-P_{k})\longrightarrow\mathrm{Ran}(I_{mn}-Q_{k}),$
i.e. a matrix $W_{k}\in\mathbb{M}_{mn,n}$ such that
$W_{k}^{*}W_{k}=I_{n}-P_{k},\qquad W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}.$
(For example, choose orthonormal bases of the two subspaces and map one to the other.)
Now define
$V_{k}:=U_{k}+W_{k}\in\mathbb{M}_{mn,n}.$
Then, using $U_{k}^{*}W_{k}=0$ (their ranges are orthogonal since $W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}$) and $W_{k}^{*}U_{k}=0$ (initial spaces are orthogonal since $W_{k}^{*}W_{k}=I-P_{k}$), we get
$V_{k}^{*}V_{k}=U_{k}^{*}U_{k}+W_{k}^{*}W_{k}=P_{k}+(I_{n}-P_{k})=I_{n}.$
Hence $V_{k}$ is an isometry in $\mathbb{M}_{mn,n}$. Moreover, since $H_{kk}^{1/2}(I_{n}-P_{k})=0$, we still have
$V_{k}H_{kk}^{1/2}=(U_{k}+W_{k})H_{kk}^{1/2}=U_{k}H_{kk}^{1/2}=R_{k}.$
Step 4: Conclude the decomposition. From $R_{k}=V_{k}H_{kk}^{1/2}$ we deduce
$R_{k}R_{k}^{*}=V_{k}H_{kk}^{1/2}\,H_{kk}^{1/2}V_{k}^{*}=V_{k}H_{kk}V_{k}^{*}.$
Summing over $k$ and using (4.2) gives
$H=\sum_{k=1}^{m}R_{k}R_{k}^{*}=\sum_{k=1}^{m}V_{k}H_{kk}V_{k}^{*},$
which is exactly (4.1). ∎
Now, we compute Fourier conjugations of two direct sums. Let $\omega=e^{2\pi i/3}$ and set
$W_{3}=\frac{1}{\sqrt{3}}\begin{pmatrix}I&I&I\\ I&\omega I&\omega^{2}I\\ I&\omega^{2}I&\omega I\end{pmatrix}=F_{3}\otimes I_{n}\in\mathbb{M}_{3n}.$
For
$D:=|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}=\operatorname{diag}(D_{1},D_{2},D_{3}),\quad\begin{cases}D_{1}=|A+B|^{2},\\ D_{2}=|B+C|^{2},\\ D_{3}=|A+C|^{2},\end{cases}$
one has
$W_{3}DW_{3}^{*}=\frac{1}{3}\begin{pmatrix}D_{1}+D_{2}+D_{3}&D_{1}+\omega^{2}D_{2}+\omega D_{3}&D_{1}+\omega D_{2}+\omega^{2}D_{3}\\ D_{1}+\omega D_{2}+\omega^{2}D_{3}&D_{1}+D_{2}+D_{3}&D_{1}+\omega^{2}D_{2}+\omega D_{3}\\ D_{1}+\omega^{2}D_{2}+\omega D_{3}&D_{1}+\omega D_{2}+\omega^{2}D_{3}&D_{1}+D_{2}+D_{3}\end{pmatrix}$. (4.3)
In particular, every diagonal block of $W_{3}DW_{3}^{*}$ equals $(D_{1}+D_{2}+D_{3})/3$.
Proof of Theorem 1.10.
Set $D=\operatorname{diag}(|A+B|^{2},|B+C|^{2},|A+C|^{2})$ and let $W_{3}$ be as above. Then $W_{3}DW_{3}^{*}\geq 0$ and, by (4.3), each diagonal block equals $(D_{1}+D_{2}+D_{3})/3$. A direct expansion shows
$D_{1}+D_{2}+D_{3}=|A+B|^{2}+|B+C|^{2}+|A+C|^{2}=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
Denote this common sum by $T$. Hence $W_{3}DW_{3}^{*}$ is a $3\times 3$ block positive matrix with constant diagonal block $T/3$. Applying (4.1) to $H=W_{3}DW_{3}^{*}$ yields isometries $V_{1},V_{2},V_{3}\in\mathbb{M}_{3n,n}$ such that
$W_{3}DW_{3}^{*}=\sum_{k=1}^{3}V_{k}\,(T/3)\,V_{k}^{*}.$
Conjugating by $W_{3}^{*}$ and setting $U_{k}=W_{3}^{*}V_{k}$ (still isometries) gives the claim. ∎
Next, let $\eta=e^{2\pi i/4}=i$ and set
$W_{4}=\frac{1}{2}\begin{pmatrix}I&I&I&I\\ I&\eta I&\eta^{2}I&\eta^{3}I\\ I&\eta^{2}I&\eta^{4}I&\eta^{6}I\\ I&\eta^{3}I&\eta^{6}I&\eta^{9}I\end{pmatrix}=F_{4}\otimes I_{n}\in\mathbb{M}_{4n},$
so that $\eta^{2}=-1$, $\eta^{3}=-i$, $\eta^{4}=1$. For
$E:=|A+B+C|^{2}\oplus|A|^{2}\oplus|B|^{2}\oplus|C|^{2}=\operatorname{diag}(E_{1},E_{2},E_{3},E_{4}),$
with $E_{1}=|A+B+C|^{2}$, $E_{2}=|A|^{2}$, $E_{3}=|B|^{2}$, $E_{4}=|C|^{2}$, we obtain
$W_{4}EW_{4}^{*}=\frac{1}{4}\bigl[M_{rs}\bigr]_{r,s=1}^{4},\qquad M_{rs}=E_{1}+\eta^{-(r-s)}E_{2}+\eta^{-2(r-s)}E_{3}+\eta^{-3(r-s)}E_{4}.$ (4.4)
In particular, every diagonal block of $W_{4}EW_{4}^{*}$ equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$.
Proof of Theorem 1.11.
Let $E=\operatorname{diag}(|A+B+C|^{2},|A|^{2},|B|^{2},|C|^{2})$ and let $W_{4}$ be the $4n\times 4n$ Fourier matrix above. Then $W_{4}EW_{4}^{*}\geq 0$ and, by (4.4), each diagonal block equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$. As in the previous proof,
$E_{1}+E_{2}+E_{3}+E_{4}=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}=|A+B|^{2}+|B+C|^{2}+|A+C|^{2}.$
Denote this sum by $S$. Thus $W_{4}EW_{4}^{*}$ is a $4\times 4$ block positive matrix with constant diagonal block $S/4$. Applying (4.1) to $H=W_{4}EW_{4}^{*}$ yields isometries $V_{1},\dots,V_{4}\in\mathbb{M}_{4n,n}$ such that
$W_{4}EW_{4}^{*}=\sum_{k=1}^{4}V_{k}\,(S/4)\,V_{k}^{*}.$
Conjugating by $W_{4}^{*}$ and setting $U_{k}=W_{4}^{*}V_{k}$ completes the proof. ∎
4.2. Clarkson–McCarthy type inequalities
Let $A,B,C\in\mathbb{M}_{n}$. Set
$Y_{1}:=\begin{pmatrix}A+B\\ B+C\\ C+A\end{pmatrix}\in\mathbb{M}_{3n,n}.$
Consider the $4\times 3$ partial Hadamard (isometry) matrix
$U=\frac{1}{2}\begin{pmatrix}1&1&1\\ 1&1&-1\\ 1&-1&1\\ 1&-1&-1\end{pmatrix},\qquad U^{*}U=I_{3},\qquad\mathcal{U}:=U\otimes I_{n}\in\mathbb{M}_{4n,3n}.$
Define
$X:=\mathcal{U}\,Y_{1}=\begin{pmatrix}A+B+C\\ B\\ A\\ -C\end{pmatrix}\in\mathbb{M}_{4n,n}.$
In particular, since |C|=|C||-C|=|C|, we have
$X^{*}X=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
Starting from the Euler-type square-sum relation and its orbit refinements, we now record several consequences for unitarily invariant norms and Schatten norms by using theorems in [8]. It suffices to notice the operator identity (1.3).
Corollary 4.6 (Unitary-invariant norms under Euler’s square-sum constraint).
[…]
For $0<p\leq 2$, both inequalities are reversed.
The preceding norm inequalities admit corresponding unitary-orbit refinements, which we state next.
Corollary 4.7 (Unitary-orbit Euler refinement).
Let $A,B,C\in\mathbb{M}_{n}$. If $p\geq 2$, then there exist unitaries $U_{1},U_{2},U_{3}\in\mathbb{M}_{n}$ such that
$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-1}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
Moreover, there exist unitaries $V_{1},V_{2},V_{3},V_{4}\in\mathbb{M}_{n}$ such that
$V_{1}|A+B+C|^{p}V_{1}^{*}+V_{2}|A|^{p}V_{2}^{*}+V_{3}|B|^{p}V_{3}^{*}+V_{4}|C|^{p}V_{4}^{*}\leq 3^{\frac{p}{2}-1}\big(|A+B|^{p}+|B+C|^{p}+|C+A|^{p}\big).$
For $0<p\leq 2$, both inequalities reverse.
Passing from general unitarily invariant norms to Schatten $p$-norms, we obtain Clarkson–McCarthy type estimates in Corollary 1.12. Using the explicit coefficient matrix $U$ above (hence $\max_{i,j}|u_{ij}|=\tfrac{1}{2}$), we obtain the mixed $\ell^{q}$$\ell^{p}$ Clarkson–McCarthy type estimate in Corollary 1.13.
4.3. A unitary-orbit Euler modulus inequality and singular value consequences
We use the following unitary subadditivity result for concave functions due to Aujla and Bourin [1, Theorem 2.1].
Lemma 4.8 (Aujla–Bourin).
Let $H,K\in\mathbb{M}_{n}$ be positive semidefinite and let $f:[0,\infty)\to[0,\infty)$ be monotone concave with $f(0)\geq 0$. Then there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$f(H+K)\leq Uf(H)U^{*}+Vf(K)V^{*}.$
Theorem 4.9.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist unitaries $U,V,W\in\mathbb{M}_{n}$ such that
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq U|A+B|U^{*}+V|B+C|V^{*}+W|C+A|W^{*}.$ (4.5)
Proof.
By (1.3),
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}=\sqrt{|A+B|^{2}+|B+C|^{2}+|C+A|^{2}}.$
Apply Lemma 4.8 with $f(t)=\sqrt{t}$ to $H=|A+B|^{2}$ and $K=|B+C|^{2}+|C+A|^{2}$ to get unitaries $U_{1},V_{1}$ with
$\sqrt{H+K}\leq U_{1}|A+B|U_{1}^{*}+V_{1}\sqrt{K}V_{1}^{*}.$
Apply Lemma 4.8 again to $|B+C|^{2}$ and $|C+A|^{2}$ to obtain unitaries $U_{2},V_{2}$ such that
$\sqrt{K}\leq U_{2}|B+C|U_{2}^{*}+V_{2}|C+A|V_{2}^{*}.$
Substitute and absorb conjugation by $V_{1}$ into $U_{2},V_{2}$ to obtain (4.5). ∎
Corollary 4.10.
[…]
Proof.
[…]
Corollary 4.11.
Let $A,B,C\in\mathbb{M}_{n}$. Then for all integers $j,k,\ell\in\{0,1,\dots,n-1\}$,
$\mu_{1+j+k+\ell}^{\downarrow}\!\Big(\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\Big)\leq\mu_{1+j}^{\downarrow}(A+B)+\mu_{1+k}^{\downarrow}(B+C)+\mu_{1+\ell}^{\downarrow}(C+A).$
Proof.
By Theorem 4.9, there exist unitaries $U,V,W\in\mathbb{M}_{n}$ such that
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq U|A+B|U^{*}+V|B+C|V^{*}+W|C+A|W^{*}.$
Set $X:=U|A+B|U^{*}$, $Y:=V|B+C|V^{*}$ and $Z:=W|C+A|W^{*}$. Then $X,Y,Z\geq 0$ and $\mu_{j}^{\downarrow}(X)=\mu_{j}^{\downarrow}(|A+B|)=\mu_{j}^{\downarrow}(A+B)$, and similarly for $Y,Z$.
We use the following singular-value inequality (see, e.g., [2]): for $P,Q\geq 0$ and $r,s\geq 0$,
$\mu_{1+r+s}^{\downarrow}(P+Q)\leq\mu_{1+r}^{\downarrow}(P)+\mu_{1+s}^{\downarrow}(Q).$ (4.6)
Applying (4.6) to $P=X$ and $Q=Y+Z$ gives
$\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z)\leq\mu_{1+j}^{\downarrow}(X)+\mu_{1+k+\ell}^{\downarrow}(Y+Z).$
Next apply (4.6) to $P=Y$ and $Q=Z$ to obtain
$\mu_{1+k+\ell}^{\downarrow}(Y+Z)\leq\mu_{1+k}^{\downarrow}(Y)+\mu_{1+\ell}^{\downarrow}(Z).$
Combining the last two inequalities yields
$\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z)\leq\mu_{1+j}^{\downarrow}(X)+\mu_{1+k}^{\downarrow}(Y)+\mu_{1+\ell}^{\downarrow}(Z).$
Finally, since
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq X+Y+Z$
and both sides are positive semidefinite, the monotonicity of singular values gives
$\mu_{1+j+k+\ell}^{\downarrow}\!\Big(\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\Big)\leq\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z),$
which implies the desired estimate. ∎
$4.$ Symmetric norms, antinorms, and Ky Fan consequences
Corollary 4.12.
Let $A,B,C\in\mathbb{M}_{n}$.
  1. (a)
[…]
  2. (b)
    For every symmetric antinorm $\|\cdot\|_{!}$ on $\mathbb{M}_{n}^{+}$ (in the sense of [3]),
    $\Big\||A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big\|_{!}\geq\big\||A+B|^{2}\big\|_{!}+\big\||B+C|^{2}\big\|_{!}+\big\||C+A|^{2}\big\|_{!}.$
Proof.
Both inequalities follow from (1.3) and the (anti)triangle inequality on $\mathbb{M}_{n}^{+}$: a symmetric norm is subadditive, while a symmetric antinorm is superadditive. ∎
Corollary 4.13.
Let $A,B,C\in\mathbb{M}_{n}$ and $m=1,\dots,n$. Denote by $\mu_{j}^{\downarrow}(\cdot)$ (resp. $\mu_{j}^{\uparrow}(\cdot)$) the singular values in nonincreasing (resp. nondecreasing) order. Then
$\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\leq\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|C+A|^{2}\big),$
$\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\geq\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|C+A|^{2}\big).$
Proof.
Apply Corollary 4.12 to Ky Fan $m$-norms and Ky Fan $m$-antinorms on $\mathbb{M}_{n}^{+}$. ∎
Acknowledgments
Teng Zhang is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No. xzy022024045).